# Chance Constrained PDE-Constrained Optimal Design Strategies Under High-Dimensional Uncertainty

### v1

@@ 3.2 Monte Carlo with control variate
The approximation of $\mathbb{E}[Q]$ and $\mathbb{V}[Q]$ in the cost functional (12) using a quadratic Taylor series expansion may lack accuracy for the design problem,particularly for design problems where the variance of the uncertain parameter $m$ is large, as shown in [31]. To enhance accuracy, the quadratic approximation $Q_{\text{QUAD}}$ is leveraged as a control variate to reduce variance in the Monte Carlo estimators for the mean and variance. This involves correcting the moments of $Q_{\text{QUAD}}$ by applying a Monte Carlo method to estimate the mean and variance of the residual between $Q$ and its quadratic approximation $Q_{\text{QUAD}}$. We can use the same $M=M_{1}=M_{2}$ random samples for the approximation of the mean and variance. For mean of control variate approximation
$\mathbb{E}[Q_{CV}]\approx\mathbb{E}[Q]=\mathbb{E}[Q_{\text{QUAD}}]+\mathbb{E}[Q-Q_{\text{QUAD}}]:=Q(\bar{m})+\frac{1}{2}tr(\mathcal{H})$ (44)
$+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-Q(\bar{m})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right).$
@@
$\mathbb{V}[Q_{CV}]:=\langle\mathcal{C}Q^{m}(\bar{m}),Q^{m}(\bar{m})\rangle+\frac{1}{4}{(tr(\mathcal{H}))}^{2}+\frac{1}{2}tr(\mathcal{H}^{2})$ (46)
$+\frac{1}{M}\sum_{i=1}^{M}\left({(Q(m_{i})-Q(\bar{m})}^{2}-{\left(\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle+\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)}^{2}\right)$
$-{\left(\frac{1}{2}tr(\mathcal{H})+\frac{1}{M}\sum_{i=1}^{M}\left(Q(m_{i})-\langle m_{i}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m_{i}-\bar{m},Q^{mm}(\bar{m})(m_{i}-\bar{m})\rangle\right)\right)}^{2}.$
The number of samples $M$ required to achieve the desired accuracy (and hence the computational cost) depends on the variance of the integrand and how closely the quadratic approximation matches the high-fidelity Monte Carlo estimation. A closer approximation reduces the required $M$, while a poorer match increases it. Notably, if the quadratic approximation is highly correlated with $Q$, fewer Monte Carlo samples are needed for effective variance reduction and accurate approximation.
3.3 Randomized trace estimator
The equations for the mean (23) and variance (27) of the design objectives evaluated using the quadratic approximation requires the trace of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}c$, which typically comes out to be a large dense matrix. The traditional approach is based on Monte Carlo method, where the Hessian action is computed on a certain number of randomly chosen directions from a Gaussian distribution [18].
Despite adhering to the lower bounds on the required number of samples for accurate trace approximations, as outlined in [48], a large number of samples is still necessary. Given the large sample size needed at each optimization iteration, this approach becomes computationally prohibitive for the design under uncertainty problem addressed in this work. Here, we implement a randomized approximated eigen-decomposition for computing the trace and trace-squared using the dominant eigenvalues $\{\lambda_{n}\}_{n=1}^{N_{\mathrm{eig}}}$ of the covariance-preconditioned Hessian ¯$\bar{\mathcal{H}}_{c}$, expressed as,
$\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n},\quad\mathrm{tr}\big{(}\bar{\mathcal{H}}_{c}^{2}\big{)}\approx\sum_{n\geq 1}^{N_{\mathrm{eig}}}\lambda_{n}^{2}.$ (47)
The above forms of trace estimators are justified by the symmetric positive definite property of the operator $\mathcal{H}_{c}$. It is important to note that the value of $N_{\mathrm{eig}}$ does not depend on the actual parameter dimension, but is dependent on the effective dimension of the parameter, e.g., [49, 50, 51, 52] it tends to be small when the covariance-preconditioned Hessian is low-rank or when its eigenvalues decay rapidly (see the numerical results of Section 5).
The eigenvalues $\lambda_{n}$ are obtained by solving the following generalized eigenvalue problem,
$\big{<}\zeta,\bar{Q}^{mm}\,\psi_{j}\big{>}=\lambda_{j}\big{<}\zeta,\mathcal{C}^{-1}\psi_{j}\big{>},\quad\forall\zeta\in\mathcal{M},\quad j=1,...,N_{\mathrm{eig}}$ (48)
where the eigenvectors $\psi_{j}$ exhibit orthonormality with $\mathcal{C}^{-1}$, $\big{<}\psi_{i},\mathcal{C}^{-1}\psi_{j}\big{>}=\delta_{ij},i,j=1,...,N_{\mathrm{eig}}$ and $\delta_{ij}$ denotes the Kronecker delta function. Using an $n$-degree finite element discretization, (48) is rewritten in the form of matrix as $\mathbf{M}\,\boldsymbol{\psi}=\lambda\,\mathbf{K}\,\boldsymbol{\psi}$ where $\mathbf{M},\mathbf{K}$ are $n$-by-$n$ symmetric matrices, and $\mathbf{K}$ is positive definite, $\boldsymbol{\psi}$ represents a vector of size $n$. For solving the resulting eigenproblem, we implement a randomized algorithm [53], discussed in Algorithm 1.
1
Input : Matrices $\mathbf{M},\mathbf{K}\in\mathbb{R}^{n\times n}$ obtained from finite element discretization, number of eigenpairs requested $N_{\mathrm{eig}}$, oversampling parameter $N_{\mathrm{p}}$
Output : Eigenpairs $(\lambda_{j},\psi_{j})$ for $j=1,\dots,N_{\mathrm{eig}}$
2
3Generate random sample and draw a Gaussian random matrix $\mathbf{\Omega}\in\mathbb{R}^{n\times(N_{\mathrm{eig}}+N_{\mathrm{p}})}$;
4
5Form the range matrix $\mathbf{Z}=\mathbf{K}^{-1}(\mathbf{M}\mathbf{\Omega})$;
6
7Perform QR factorization $\mathbf{Z}=\mathbf{Q}\mathbf{R}$ such that $\mathbf{Q}^{\top}\mathbf{M}\mathbf{Q}=\mathbf{I}_{N_{\mathrm{eig}}+N_{\mathrm{p}}}$;
8
9Compute the reduced matrix $\mathbf{Q}^{\top}\mathbf{K}\mathbf{Q}$;
10
11Compute the eigendecomposition $\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\dots)$;
12
13Obtain the eigenpairs of the original problem:
$\lambda_{j}=\boldsymbol{\lambda}[j,j],\quad\psi_{j}=\mathbf{Q}\mathbf{U}[:,j],\quad j=1,\dots,N_{\mathrm{eig}}.$
Algorithm 1 Double-Pass Randomized Eigenvalue Solver
@@ 4 Gradient-based optimization
4.1 Smooth approximation of the discontinuous function
The indicator function utilized in the chance constraint exhibits discontinuity at $f(m,d)=0$. To employ the gradient-based INCG, this discontinuous indicator function must be approximated with a smooth, continuous counterpart. One approach involves employing a logistic function as a smooth approximation of the indicator function [25] such that,
$\mathbb{I}_{[0,\infty]}(x)\approx l_{\omega}(x)=\frac{1}{1+e^{-2\omega x}},$ (49)
where a larger $\omega$ corresponds to a sharper transition at $x=0$ as shown in Figure 3. This smooth approximation tackles the non-differentiability of the chance constraint and makes the cost functional differentiable with respect to the design variable.
Figure 3: Approximation of the discontinuous indicator function $\mathbb{I}_{[0,\infty]}(x)$ with a smooth logistic function as given in (49) and its convergence with increasing $\omega$
@@ 4.2 Penalty Method
To impose the inequality constraint (15), a quadratic penalty method [55] is applied. The quadratic penalty function is defined as,
$S_{\gamma}(x)=\frac{\gamma}{2}{(max{\{0,x\}})}^{2}$ (50)
where $\gamma>0$ is a constant corresponding to the weight of the penalty. Using the penalty method, the chance-constrained problem can be rewritten as
$\underset{d\in[0,1]}{\text{min}}\mathcal{J}(d)+\mathcal{S}_{\gamma}(\mathbb{E}[l_{\omega}(f)]-\alpha_{c})$ (51)
This transformation simplifies the inequality-constrained problem by converting it into an unconstrained form.
4.3 Design gradients via adjoint
The Lagrangian formalism is employed to derive the gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter, referred to as the d-gradient. This gradient provides critical information to the INCG optimizer, in order to facilitate gradient-based optimization processes. By substituting the trace estimators (47) into (28) and converting it into unconstrained problem using quadratic penalty (50) and smooth approximation (49), we have
@@ 4.3 Design gradients via adjoint
$\mathcal{J}_{\mathrm{QUAD}}(d)$ $=\bigg{(}\bar{Q}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\lambda_{j}^{Q}\bigg{)}+\beta_{V}\bigg{(}\big{<}\bar{Q}^{m},\mathcal{C}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\left({\lambda_{j}^{Q}}\right)^{2}\bigg{)}+\beta_{R}\,R(d)$ (52)
$+S_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
To evaluate the $d$-gradient, we consider a quadratic meta-Lagrangian functional
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\left(\psi_{j}^{Q}\right)^{*},(\bar{Q}^{mm}-{\lambda_{j}^{Q}}\mathcal{C}^{-1})\psi_{j}^{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\left(\lambda_{j}^{Q}\right)^{*}(\langle{\psi_{j}}^{Q},\mathcal{C}^{-1}{\psi_{j}}^{Q}\rangle-1)+\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle{\hat{\mathbf{v}}_{j}}^{{Q}^{*}},\partial_{\mathbf{vu}}\bar{r}_{Q}\,{\hat{\mathbf{u}}_{j}}^{Q}+\partial_{\mathbf{v}m}\bar{r}_{Q}\,{\psi_{j}}^{Q}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\psi_{j}^{f}\right)^{*},(\bar{f}^{mm}-{\lambda_{j}^{f}}\mathcal{C}^{-1})\psi_{j}^{f}\rangle$
$+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\left(\lambda_{j}^{f}\right)^{*}(\langle{\psi_{j}}^{f},\mathcal{C}^{-1}{\psi_{j}}^{f}\rangle-1)+\sum^{N_{\mathrm{eig}}^{f}}_{j=1}\langle\left(\hat{\mathbf{v}}_{j}^{f}\right)^{*},\partial_{\mathbf{vu}}\bar{r}_{f}\,{\hat{\mathbf{u}}_{j}}^{f}+\partial_{\mathbf{v}m}\bar{r}_{f}\,{\psi_{j}}^{f}\rangle$
where $(\cdot^{*})$ represents the adjoint variables, $(\hat{\cdot})$ denotes the incremental variables. The terms on the right hand side of (53), correspond to the following: the quadratic approximation of the design objective (28), the state (32) and adjoint (33) problems for design objective $Q$, the generalized eigenvalue problems (48) for design objective $Q$, the orthonormality constraint of the eigenvectors for design objective $Q$, the incremental state problem (36) of design objective $Q$ for eigenvector directions, the incremental adjoint problem (37) of design objective $Q$ for the directions of the eigenvectors, the adjoint problem (39) for chance constraint $f$, the generalized eigenvalue problems (48) for chance constraint $f$, the orthonormality constraint of the eigenvectors for chance constraint $f$, the incremental state problem (41) of chance constraint $f$ and the incremental adjoint problem (42) of chance constraint for the directions of the eigenvectors.
To compute the design gradient, we need to calculate five sets of functional derivatives to solve for the corresponding ten sets of adjoint variables
$\left(\mathbf{u}^{Q}\right)^{*},\left(\mathbf{u}^{f}\right)^{*},\left(\mathbf{v}^{Q}\right)^{*},\left(\mathbf{v}^{f}\right)^{*},\left(\psi_{j}^{Q}\right)^{*},\left(\psi_{j}^{f}\right)^{*},\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*},\left(\hat{\mathbf{u}^{f}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{Q}_{j}}\right)^{*},\left(\hat{\mathbf{v}^{f}_{j}}\right)^{*}\}$ in (53), while the adjoint variable set for the eigenvalue set $\{{\lambda_{j}^{Q}}^{*},{\lambda_{j}^{f}}^{*}\}$ is not required due to its independency of the design parameter $d$.
Setting the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}\}$ to zero isolates the terms corresponding to the quadratic approximation of the design objective and the orthonormality constraint of the eigenvectors, resulting in:
$\{{\psi^{Q}_{j}}^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda_{j}^{Q}\bigg{)}\,\{\psi_{j}^{Q}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (54)
Setting the first variations of (53) to zero with respect to $\mathbf{v}^{Q}_{j}$ and $\mathbf{u}^{Q}_{j}$ results in the incremental state problem:
$\small\textit{Find $\hat{\mathbf{u}}^{Q}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (55)
and the incremental adjoint problem:
$\footnotesize\begin{split}\textit{Find $\hat{\mathbf{v}}^{Q}\in\mathcal{V}$, such that}&\quad\\ \big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}_{j}\right)^{*}\big{>}=-\big{<}\tilde{\mathbf{u}},\partial_{\mathbf{uu}}\bar{r}^{Q}\,{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{uu}}\bar{Q}\,&{\left(\hat{\mathbf{u}}^{Q}_{j}\right)}^{*}+\partial_{\mathbf{u}m}\bar{r}^{Q}\,\left(\psi^{Q}_{j}\right)^{*}\big{>},\quad\forall\tilde{\mathbf{u}}\in\mathcal{U},\end{split}$ (56)
similar to (36) and (37), respectively. In combination with (54), we can formulate the adjoint variables for the incremental state and adjoint problems individually for each eigenvector as
$\{\left(\hat{\mathbf{u}}\right)^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{u}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q},$ (57)
$\{\hat{\mathbf{v}}^{*}_{j}\}$ $=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{Q}_{j}\bigg{)}\,\{\hat{\mathbf{v}}^{Q}_{j}\},\quad j=1,\cdots,N_{\mathrm{eig}}^{Q}.$ (58)
By vanishing the first variation of the Lagrangian functional (53) with respect to the adjoint or the state, we obtain the linear state problem: Find ${\mathbf{u}^{Q}}^{*}\in\mathcal{U}$, such that
$\langle\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{Q}\,\left(\mathbf{u}^{Q}\right)^{*}\rangle=$ $-2\beta_{V}\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (59)
$-\sum^{N_{\mathrm{eig}}^{Q}}_{j=1}\langle\tilde{\mathbf{v}},\partial_{\mathbf{vuu}}\bar{r}^{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{vu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,\quad\forall\tilde{\mathbf{v}}\in\mathcal{V},$
and linear adjoint problem: Find $\left(\mathbf{v}^{Q}\right)^{*}\in\mathcal{V}$, such that
$\langle\tilde{\mathbf{u}},\partial_{\mathbf{uv}}\bar{r}^{Q}\,\left(\mathbf{v}^{Q}\right)^{*}\rangle=$ $-\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}}\bar{Q}\rangle-2\beta_{V}\langle\tilde{\mathbf{u}},\partial_{\mathbf{u}m}\bar{r}^{Q}\,(\mathcal{C}\partial_{m}\bar{r}^{Q})\rangle$ (60)
$+\partial_{\mathbf{uuu}}\bar{Q}\,\hat{\mathbf{u}}^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}+\partial_{\mathbf{uuu}m}\bar{r}^{Q}\,\psi^{Q}_{j}\,\left(\hat{\mathbf{u}}^{Q}_{j}\right)^{*}\rangle,$
$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
By vanishing the first variation of (53) with respect to the eigenvalue set $\{\lambda_{j}^{f}\}$, we get
$\{\left(\psi^{f}_{j}\right)^{*}\}=\bigg{(}\frac{1}{2}+\beta_{V}\,\lambda^{f}_{j}\bigg{)}\,\{\psi^{f}_{j}\},\quad j=1,\cdots,N^{f}_{\mathrm{eig}}.$ (61)
Equating to zero of the first variations of (53) with respect to $\mathbf{v}_{j}^{f}$ and $\mathbf{u}_{j}^{f}$ results in the incremental state problem:
$\textit{Find $\hat{\mathbf{u}}^{f}\in\mathcal{U}$, such that}\quad\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{vu}}\bar{r}^{f}\,\mathbf{u}^{f*}_{j}\big{>}=-\big{<}\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}\,\psi^{{f}^{*}}_{j}\big{>},\quad\forall\tilde{\mathbf{v}}\in\mathcal{V}.$ (62)
and linear adjoint problem:
Find $\left(\mathbf{v}^{f}\right)^{*}\in\mathcal{V}$, such that (63)
$-\langle\tilde{\mathbf{v}},\partial_{\mathbf{v}m}\bar{r}^{f}m^{f}\rangle$
$\quad\forall\tilde{\mathbf{u}}\in\mathcal{U}.$
Finally, with driving all the adjoint variables, the $d$-gradient is derived as
@@
The methodology for calculating the design gradient is detailed in Algorithm 2. The computational cost required to evaluate the quadratic approximation of the cost functional, denoted as $\mathcal{J}_{\mathrm{QUAD}}(d)$ in (52), includes solving one potentially nonlinear state partial differential equation (PDE), one linear adjoint problem, and $22(N_{\mathrm{eig}}^{Q}+N_{o})+2(N_{\mathrm{eig}}^{f}+N_{o})+$ 2 linear PDEs, as prescribed by the double-pass randomized algorithm presented in Algorithm 1. Furthermore, obtaining the $d$-gradient involves solving $2(N_{\mathrm{eig}}^{Q}+N_{\mathrm{eig}}^{f}+3)$ additional linear PDEs.
1
Input : Design objective (12) and weak form of the PDF (7).
Output : Gradient of the quadratic approximation of the design objective $\mathcal{J}_{\mathrm{QUAD}}(d)$ with respect to the design parameter $z$.
2
3Solve for $m$-gradient (34);
4
5Solve for $m$-Hessian (38);
6
7Solve for $m_{f}$-gradient ;
8
9Solve for $m_{f}$-Hessian ;
10
11Solve for generalized eigenpairs with Algorithm 1;
12
13Compute $\mathcal{J}_{\mathrm{QUAD}}$ by (52);
14
15Solve the linear state problem (59);
16
17Solve the linear adjoint problem (63);
18
19Obtain the design gradient (64)
Algorithm 2 Analytical design gradient

### v2

@@ 3.2 Monte Carlo with control variate
The approximation of $\mathbb{E}[Q]$ and $\mathbb{V}[Q]$ in the cost functional (7) using a quadratic Taylor series expansion may lack accuracy for the design problem, particularly for design problems where the variance of the uncertain parameter $m$ is large, as shown in [24]. To enhance accuracy, the $Q_{\text{QUAD}}$ is leveraged as a control variate to reduce variance in the sampling estimations of the mean and variance. In particular, A Monte Carlo method is employed to correct the moments of $Q_{\text{QUAD}}$ by computing the mean and variance of the residual between $Q$ and its quadratic approximation $Q_{\text{QUAD}}$. For mean of control variate approximation
$\mathbb{E}[Q_{CV}]\approx\mathbb{E}[Q]=\mathbb{E}[Q_{\text{QUAD}}]+\mathbb{E}[Q-Q_{\text{QUAD}}]:=Q(\bar{m})+\frac{1}{2}tr(\bar{\mathcal{H}}_{k})$ (24)
$+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left(Q(m^{(i)})-Q(\bar{m})-\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right).$
@@
$\mathbb{V}[Q_{CV}]:=\langle\mathcal{K}Q^{m}(\bar{m}),Q^{m}(\bar{m})\rangle+\frac{1}{4}{(tr({\bar{\mathcal{H}}_{k}}))}^{2}+\frac{1}{2}tr({\bar{\mathcal{H}}_{k}}^{2})$ (26)
$+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left({(Q(m^{(i)})-Q(\bar{m})}^{2}-{\left(\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle+\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right)}^{2}\right)$
$-{\left(\frac{1}{2}tr({\bar{\mathcal{H}}_{k}})+\frac{1}{N_{mc}}\sum_{i=1}^{N_{mc}}\left(Q(m^{(i)})-\langle m^{(i)}-\bar{m},Q^{m}(\bar{m})\rangle-\frac{1}{2}\langle m^{(i)}-\bar{m},Q^{mm}(\bar{m})(m^{(i)}-\bar{m})\rangle\right)\right)}^{2}.$
The number of samples $N_{mc}$ required to achieve the desired accuracy (and hence the computational cost) depends on the variance of the integrand and how closely the quadratic approximation matches the high-fidelity Monte Carlo estimation. Notably, if the quadratic approximation is highly correlated with $Q$, fewer Monte Carlo samples are needed for effective variance reduction and accurate approximation.
@@ 4 Gradient-based optimization
4.1 Continuous approximation of the discontinuous function
The indicator function utilized in the chance constraint is discontinuous at $f(m,d)=0$. To employ the gradient-based INCG, this discontinuous indicator function must be approximated with a continuous function. A common approach involves using a sigmoid function to approximate the indicator function, as proposed by [21], given by
$\chi_{[0,\infty)}(x)\approx l_{\omega}(x)=\frac{1}{1+e^{-2\omega x}},$ (27)
where the parameter $\omega$ controls the sharpness of the transition around $x=0$, as illustrated in Figure 3. A higher value of $\omega$ results in a steeper transition. This smooth approximation addresses the non-differentiability of the chance constraint, enabling the cost functional to be differentiable with respect to the design variable. The value of $\omega$ is gradually increased through a continuation scheme discussed in Algorithm 2. This gradual scaling of $\omega$ helps in avoiding local optima and prevents numerical instabilities.
Figure 3: Approximation of the discontinuous indicator function $\chi_{[0,\infty)}(x)$ with a sigmoid function as given in (27) and its convergence with increasing $\omega$
@@ 4.2 Penalty Method
In order to address the inequality constraint in (10), we convert the original chance-constrained optimization problem into the following unconstrained problem,
$\underset{d\in[0,1]}{\text{min}}\mathcal{J}(d)+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f)]-\alpha_{c}).$ (28)
Here, $\tau_{\gamma}$ is a quadratic penalty method as outlined in [47].The quadratic penalty function is given by,
$\tau_{\gamma}(x)=\frac{\gamma}{2}{(max{\{0,x\}})}^{2},$ (29)
where $\gamma>0$ is a constant which determines the penalty weight. In this manner, the inequality constraint is replaced with a penalization term, that results in the transformation of the constrained problem into an unconstrained form. As $\gamma$ increases as per the continuation scheme in Algorithm 2, the penalized solution approaches the feasible region more strictly, making the unconstrained optimization problem effectively mimic the constrained form.
4.3 Design gradient
By substituting the trace estimators (23) into (18) and converting it into unconstrained problem using quadratic penalty (29) and continuous approximation (27), we have
@@ 4.3 Design gradient
$\mathcal{J}_{\mathrm{QUAD}}(d)$ $=\bigg{(}\bar{Q}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\lambda_{j}^{Q}\bigg{)}+\beta_{V}\bigg{(}\big{<}\bar{Q}^{m},\mathcal{K}\,\bar{Q}^{m}\big{>}+\frac{1}{2}\sum_{j\geq 1}^{N_{\mathrm{eig}}}\left({\lambda_{j}^{Q}}\right)^{2}\bigg{)}+\beta_{RG}\,R_{p}(d)$ (30)
$+\tau_{\gamma}(\mathbb{E}[l_{\omega}(f_{\text{QUAD}}(m,d))]-\alpha_{c})$
Implementing the INCG optimizer requires computing the gradient of the quadratic approximation of the design objective, $\mathcal{J}_{\mathrm{QUAD}}(d)$, with respect to the design parameter, referred to as the $d$-gradient. The Lagrangian formalism of the approximated cost, along with variational calculus principles for obtaining the necessary Lagrange multipliers, and an algorithm for computing the $d$-gradient are presented in Section B.
@@
The computational cost required to evaluate the quadratic approximation of the cost functional, denoted as $\mathcal{J}_{\mathrm{QUAD}}(d)$ in (30), includes solving one potentially nonlinear state PDE, one linear adjoint problem, and $22(N_{\mathrm{eig}}^{Q}+N_{o})+2(N_{\mathrm{eig}}^{f}+N_{o})+$ 2 linear PDEs, as prescribed by the double-pass randomized algorithm (see [24, 14]). Furthermore, obtaining the $d$-gradient involves solving $32(N_{\mathrm{eig}}^{Q}+N_{\mathrm{eig}}^{f})+$ 3 additional linear PDEs. The design gradient is needed for obtaining the optimal design through a continuation scheme as discussed in Algorithm 2.